# Artworks Reimagined: Exploring Human-AI Co-Creation through Body Prompting

### v1

@@ 6. Conclusion
Generative AI opens new opportunities and alternative ways to engage diverse audiences in image generation. Embodied interaction for co-creating images with generative AI is an intriguing possibility, providing a novel, fun, and highly engaging experience. Our study explored embodied interaction (body prompting) as a way to interact with image generation AI in public spaces. As demonstrated by our study, such interaction acts as a meaningful connector between historical artistic practices, such as the act of posing for a portrait, and new creative processes involving AI technology. As observed in our installation, generative AI provides a means for effectively fostering personal expression and collaborative co-creation. Our recommendations serve as guidance for future studies and the design of installations involving generative AI in the context of GLAM institutions and other public settings.
Acknowledgements.
This work is supported by the Natural Sciences and Engineering Research Council of Canada (NSERC). We thank the research assistants Aleksandra Hämäläinen, Saila Karjalainen, and Petri Salo for their contributions to this research project.

### v2

@@ 7. Conclusion
Generative AI opens new opportunities and alternative ways to engage audiences in image generation with generative AI. Embodied interaction for co-creating images with generative AI is an intriguing possibility, providing a novel, fun, and highly engaging experience. Our study explored body prompting as a way to interact with image generation AI via an art installation in a public event setting. As demonstrated by our study, such interaction acts as a meaningful connector between historical artistic practices, such as the act of posing for a portrait, and new creative processes involving AI technology. As observed in our installation, generative AI provides a means for effectively fostering personal expression and collaborative co-creation. Our recommendations serve as guidance for future studies and the design of installations involving generative AI in the context of GLAM institutions and other public settings.
Acknowledgements
This work was partially supported by the Natural Sciences and Engineering Research Council of Canada (NSERC). We thank reviewers for their valuable input. We are grateful to the research assistants Aleksandra Hämäläinen, Saila Karjalainen, and Petri Salo for their contributions to this research project.